Ahmer walks 300m east and 80m north 280m east find the displacement and distance?
1
Expert's answer
2020-03-05T09:37:32-0500

Distance:


"d=300+80+280=660\ m"

Displacement:


"s=\sqrt{(300+280)^2+80^2}=585\ m"

Direction:


"\theta=\arctan{\frac{80}{580}}=7.85\degree\text{N of E}"